OpenStudy (anonymous):

If f(x) = 8 – 3|x + 2|, find f(–6) A. 4 B. –4 C. 20 D. –16

undeadknight26 (undeadknight26):

f(-6) = 8 - 3I-6 +2I

undeadknight26 (undeadknight26):

now solve to get your answer :D

OpenStudy (anonymous):

just plug in -6!!!!!

undeadknight26 (undeadknight26):

don't forget when you do -6 + 2 = -4 you ahve to turn it into positive so:8 - 3 x 4.

OpenStudy (anonymous):

So is it 5x4? which is 20

undeadknight26 (undeadknight26):

oh i thought it was 8 - 12 = -4...

undeadknight26 (undeadknight26):

now im not 100% sure im right -_-...

OpenStudy (anonymous):

Well thank you either way I couldn't figure out what i was doing thanks!

undeadknight26 (undeadknight26):

@surjithayer is typing...

OpenStudy (anonymous):

\[\left| x+2 \right|=-\left( x+2 \right) ,if x+2 <0 or x<-2,-6<-2\] f(x)=8+3(x+2)=8+3x+6=3x+14 f(-6)=3*-6+14=-18+14=-4